Shift Summary for day shift, 24/jul/99

Ashkan_Alavi-Harati (alavi@fnal.gov)
Sat, 24 Jul 1999 16:15:04 -0500

Shift Summary for day shift, 24/jul/99

On Shift: Albert Erwin, Ashkan Alavi

Hours of beam delivered at all intensities: 4.0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 3.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:accelerator problems

Main Injector vacuum problem, TeV quench, dart fresh starts.

========================================================================
Runs ended this shift:

14010 BEAM_832 6.3M evnets
14011 BEAM_832 380K evnets

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift:none

========================================================================
Overview of shift:

E832 data taking. The sem was about 2.5-3.3E12 (request 4.2E12) and BOA
factor about 2 most of the time. Poor beam quality and interrupted by
various TeV problems.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14010 120 07:00 Take over the shift. We have good beam on the
target but SEM is only 3.2 E12.

161 08:05 BOA is about 1.8, getting worse.
171 08:16 Called the control room to complain about the BOA
factor (2.2 now).
200 09:00 DYC full, KFREND fixes.
223 09:24 BOA still high. Contacted Rick Ford.
09:50 Rick Ford got in touch with the operators. They
are working on the beam quality. If the BOA does
not improve in half and hour, he will bug the TeV
people.
255 10:07 No beam, Vacuum control problems in MI.
10:49 Estimate 45 minutes.
325 11:40 Stop the run. Do drt_fresh_start.

10411 2 12:02 Start the run. No beam yet, no updates.
47 13:06 Beam returns.
13:20 DYC full, KFREND fixes.
13:40 Pipeline error. Can not open the pipeline window on
any screens!
13:45 Stop the run to drt_fresh_start.

10412 1 14:00 Start the new run. Pipeline problem seems to be
fixed.
7 14:10 TeV quench at E4, no beam.

========================================================================
Follow up Notes on Previous Shift

The frequency on CsI 961 with comparator errors has increased from 41 to
125 in about 5 hours. Experts are aware of the problem.

========================================================================
Items to Follow up on Next Shift

Keep an eye on the problematic comparator by PAW > pipe#qielist

One of the DCOR2 power supplies is a bit higher than the nominal values.
Ignore the error, experts (the shift crew!) know about it.